# Accelerating Large Language Model Pretraining via LFR Pedagogy: Learn, Focus, and Review

### v1

@@ 4 LFR Training Methodology
    Phase 1: Learn for 1 epoch ($p_{1}=1$).
@@
    Phase 2: Focus on 50% of the data for 1 epoch ($s_{1}=50,p_{2}=1$).
@@
    Phase 3: Review the entire dataset for another epoch ($p_{3}=1$).
  4. 4.
    Phase 4: Focus on 30% of the data for 5 epochs ($reps=2,s_{2}=70,p_{4}=1$).
We choose this configuration based on our limited pretraining budget and the profiling in Section 3.2 which revealed that $\sim$25of data samples are forgotten at least once during training. From our observations in Figure 1, we notice that the data samples that are forgotten are forgotten multiple times. On average, they had to be presented to the model 4 times for them to be learned. Therefore, we use the first three training phases to identify these samples, and then spend 5 epochs focusing on them. To be on the conservative side, we focus on 30% of the samples in the last 5 epochs, to ensure that the model retains these difficult samples in its long-term memory. Furthermore, we ensure that all data samples are reviewed through Phases 1 and 3 of our approach. Finally, we test other values for $p_{1}$, $s_{1}$, $p_{2}$, $p_{3}$, and $reps$ and evaluate the sensitivity of LFR to these hyperparameters in Section 5.9.
Algorithm 1 LFR Training Methodology
Training dataset $D$, model $M$ with initial parameters $\theta_{0}$, hyperparameters $p_{1}$, $s_{1}$, $p_{2}$, $p_{3}$, and $reps$.
Minimization of Equation 3.
$PPLs,\theta_{p_{1}}\leftarrow\textbf{Learn}(\theta_{0},D,p_{1})$
for $r=1,2,\ldots,reps$ do
     $Sort(PPLs,D),S_{1}\leftarrow(100-s_{1})\%D$
     $\theta_{p_{2}}\leftarrow\textbf{Focus}(\theta_{p_{1}},S_{1},p_{2})$
     $PPLs,\theta_{p_{3}}\leftarrow\textbf{Review}(\theta_{p_{2}},D,p_{3})$
end for
$Return\ \theta$
Our training strategy is simple, intuitive and human-like. It gives the model an opportunity to learn from all of the data, prioritize and relearn forgotten data points, and repeatedly review harder data samples more frequently than they would have been using random sampling. While the clustering-based techniques Tirumala et al. (2023); Abbas et al. (2023); Kaddour (2023) presented in Section 2 allow for accelerated training, they are not designed to suit the multi-descent training dynamics observed in Section 4 and require pretrained model embeddings to calculate distance metrics. By dropping data points with high similarity (within the same cluster), the model is not given a chance to repeatedly learn harder data points and observe the multi-descent behaviour in Figure 1. Furthermore, prior methods including perplexity-based pruning methods Marion et al. (2023) are static. Sections 5.8 and A.2 characterize the data blocks found easy and hard by the LLM, and demonstrate why static, clustering-based data selection methods achieve poor downstream task performance. Lastly, our approach does not require any pretrained models to obtain embeddings. Our focused training strategy allows the model to absorb harder information (data blocks with higher perplexity) faster, by presenting them more number of times. We evaluate the performance of LFR on six downstream tasks in Section 5.
5 Evaluation
In this section, we present a comprehensive evaluation of our proposed data selection method, LFR. We pretrain GPT-2 models Radford et al. (2019) of various sizes (124M, 345M, 774M, 1.5B) on the OpenWebText dataset ope from scratch using 4 AMD MI210 GPUs (345M-1.5B) and 4 AMD MI100 GPUs (124M). Our pretraining experiments utilize a fully distributed data parallel (DDP) approach.
We also evaluate our pretrained models on downstream tasks from different domains in Sections 5.3, 5.4, and 5.5 by finetuning it with the Huggingface transformers library Wolf et al. (2020). Section 5.8 presents an analysis on the data blocks dropped and retained by LFR in different phases of training across the different model sizes. Finally, we conduct a sensitivity study in Section 5.9, where we test LFR with different hyperparameters.
5.1 Models
The GPT-2 models implement a decoder-only transformer architecture introduced in Radford et al. (2019). The model configurations are shown in Table 6 in Section A. The models were originally pretrained on the closed-source WebText dataset Radford et al. (2019) for 800k iterations. Researchers developed the OpenWebText dataset as an open-source clone of the WebText dataset by following the same data scraping method described by OpenAI in their GPT-2 release Radford et al. (2019). LFR utilizes the OpenWebText dataset for pretraining for a total of 40k iterations, i.e., 20x fewer iterations than the OpenAI release. We build upon the GPT-2 training implementation provided by Nan and train our models with the four phases described in Section 4.
@@ 5.1 Models
5.2 Baselines
  1. 1.
    OpenAI checkpoints: we compare our models pretrained through LFR for 40k iterations with the same GPT-2 architectures pretrained by OpenAI for 800k iterations. We use the same batch size as OpenAI (512) as shown in Section A.1. Due to insufficient GPU memory, we adjust the batch size by using a higher gradient accumulation. The checkpoints are sourced from Huggingface [^1] https://huggingface.co/openai-community.
  2. 2.
    Random Sampling: while the previous baseline ensures that we compare with industry-standard models, we also develop and compare LFR against the same GPT-2 architectures pretrained for 40k iterations on the OpenWebText dataset using random sampling. This baseline enables LFR to produce higher quality models than those obtained through traditional autoregressive modeling when using the exact same dataset and iteration count. Note that we are only able to produce and pretrain this baseline for the 124M, 345M, and 774M models due to limited compute resources.
  3. 3.
    Prior works: We pretrain the GPT-2 124M parameter models on the dataset generated through DSIR Xie et al. (2023) and MiniPile Kaddour (2023). More details of these baselines can be found in Section 5.7.
5.3 Language Modeling
To evaluate the performance of a retrained language model, we finetune on a variety of datasets and evaluate the performance in terms of $PPL$. Specifically, we consider WikiText-2 wik , WikiText-103 wik , LAMBADA Paperno et al. (2016), and One Billion Word Benchmark (1BW) Chelba et al. (2014) and show the results of our method and baselines in Table 2.
Compared with the OpenAI baseline, we observe large improvements with a decrease in $PPL$ ranging from 0.82-3.51 (5.3%-17.7%) for the WikiText-2 dataset across model sizes. The improvements are more significant for the larger WikiText-103 dataset. We observe a $PPL$ reduction by 3.58-9.09 (20.7%-28.8%) across model sizes. In the case of the LAMBADA dataset which measures the LLM’s ability to capture long-range dependencies in text, LFR’s focused learning approach outperforms the OpenAI checkpoints by lowering perplexity by 0.56-1.39 (3.9%-8.9%). Finally, for the 1BW benchmark, which measures progress in statistical language modeling, LFR outperforms the OpenAI models by a large reduction in $PPL$ ranging from 2.69-6.91 (10.1%-17.6%).
Based on the results in Table 2, we find that the $PPL$ reduction obtained by LFR increase as the dataset size increases (from WikiText-2 to 1BW). Also, smaller models show a larger $PPL$ reduction by using LFR than larger models. On average, using our approach, perplexity was reduced by 4.92, 3.26, 2.17, and 1.40 for the GPT-2 124M, 345M, 774M, and 1.5B models, respectively.
5.4 Translation
The WMT-14 English-French dataset Artetxe et al. (2018); wmt consists of pairs of sentences in English and French. While, non-English content is filtered out in the OpenWebText dataset (similarly as in WebText), the GPT-2 1.5B parameter model pretrained through LFR can achieve a BLEU score of 11.8 on the WMT-14 English-French test set. In contrast, the GPT-2 1.5B OpenAI baseline achieves a lower BLEU score of 11.5. Note that the GPT-2 models are not designed for translation tasks, but can still outperform unsupervised machine translation approaches Artetxe et al. (2018).
5.5 Question-Answering
Question-answering is an important domain for LLMs and is a potential test for the factual information aggregated in the parameters during training. Similarly to the GPT-2 paper, we test our pretrained models on the Natural Questions dataset Kwiatkowski et al. (2019) by augmenting our model with a linear layer to compute the start and end logits of the model output. The GPT-2 1.5B parameter baseline OpenAI model achieves 4.1% accuracy on the questions while LFR achieves a slightly higher accuracy of 4.61%. Considering that model capacity plays a major role in improving accuracy for question-answering tasks, we believe that applying LFR’s training methodology to larger models will yield even better results.
5.6 World Knowledge and Problem Solving
We evaluated LFR on the MMLU benchmark Hendrycks et al. (2021) to test the model’s ability to solve multiple-choice questions across a wide range of 57 subjects, including mathematics, US history, computer science, law, and social sciences. We compared the GPT-2 1.5B model, trained with LFR, against the OpenAI baseline in a five-shot setting. This resulted in an average accuracy improvement of 0.2% across all 57 subjects. Table 3 displays the accuracy results.
Subject 1.5B-OpenAI 1.5B-LFR STEM 24.5% 26.1% Humanities 24.8% 27.2% Social Sciences 24% 23.8% Other (business, health, misc.) 27.8% 25.1% Average (across 57 subjects) 25.3% 25.5%
Table 3: Accuracy results for the MMLU benchmark for LFR and the OpenAI Baseline for the 1.5B parameter model.
@@ 5.7 Comparison with Prior Work
In this section, we use the 124M parameter model to compare LFR with DSIR Xie et al. (2023) and MiniPile Kaddour (2023), which propose importance resampling and cluster-based data sampling respectively. In the case of DSIR, we use the OpenWebText dataset as the raw dataset and set the target dataset as WikiText-103 for selecting a subset of 1M documents ($\sim$6GB). It took a total of 2.7 hours for the data selection process on 8 AMD EPYC 7V13 CPU cores operating at 2.45GHz with 64GB RAM. For the MiniPile baseline, we pretrain the GPT-2 124M model for 40k iterations with the dataset available on Huggingface (min, ). To ensure a fair comparison, we finetune each model using 4 AMD MI100 GPUs on the same downstream tasks used in Section 5.3. Table 4 displays the downstream task perplexity. LFR significantly outperforms prior work on data selection and allows the pretrained LLM to generalize better to new finetuning objectives, by focusing and revising different parts of the training corpus at regular intervals.
5.8 Analysis on Dropped and Retained Data Blocks
  1. 1.
  2. 2.
  3. 3.
    Are the dropped data blocks similar across model sizes?
  4. 4.
To answer the first question, we printed out the texts dropped and retained at different training phases. Tables 12 and 14 show text blocks dropped in Phases 2 and 4 by the 345M and 124M parameter models, while Tables 13 and 15 show data blocks retained. By reading through the texts, we notice that the model first learned conversations and personal anecdotes, before being able to retain factual information. We provide a more detailed analysis of the learning process in Section A.2.
To summarize, data block importance varies across training time, and across model sizes. Therefore, static data selection techniques Tirumala et al. (2023); Abbas et al. (2023); Kaddour (2023); Xie et al. (2023) which select a fixed subset to train for the entire training duration for all model architectures do not adapt to the changing training dynamics of LLMs. Based on our analysis in Figure 3 and 2, different data blocks are found difficult by models of different capacities at different training instants, driving the need for dynamic data selection methods like LFR. We detail further analysis on the selected and discarded data blocks and demonstrate how models initially focus on learning conversational and anecdotal data, before proceeding to learn factual data in Appendix A.2.
5.9 Sensitivity Study
In this section, our goal is to understand the effects of more aggressive focus, revision, and learning strategies than the 4-phase training strategy presented in Section 4. Here, we vary the values of hyperparameters $p_{1},s_{1},p_{2},p_{3},$ and $reps$ and study the effects on the downstream task perplexity. Specifically, we aim to answer the following two questions:
  1. 1.
    What is the impact of removing Phase 3?
  2. 2.
    What is the impact of the degree of pruning in Phases 2 and 4?
To answer the first question, we pretrain a 124M parameter GPT-2 model without the reintroduction of data blocks in Phase 3, and use the reduced subset from Phase 2 for the rest of the training. Then, we finetune for downstream language modeling tasks similarly to in Section 5.3 and compared the perplexities using LFR in Table 5. This training strategy which removes Phase 3, is labeled as no-reintro. Next, to answer the second question, we pretrain a 124M parameter GPT-2 model using LFR but increase the degree of pruning in Phase 2 from 50% to 70%, i.e., reduce the training subset to 30% of the original size. This aggressive training strategy is labeled as aggr-2.
@@ 8 Limitations and Ethical Considerations
    LFR is evaluated on models up to 1.5B parameters using the OpenWebText dataset, constrained by our compute resources. We encourage researchers to validate such focused learning approaches for different model families, and domains.
@@ A.1 Experiment Details
(Radford et al., 2019) demonstrates that these datasets are out-of-domain for the WebText dataset using Bloom filters containing 8-grams of the training tokens. While we have not conducted a similar study for the OpenWebText dataset, since both datasets are scraped and pre-processed using a similar method, we estimate that the percentages of 8 grams overlap with the training set will be similar.
Models: Table 6 describes the different GPT-2 configurations used in LFR.
@@
Training Runtime Table 8 describes the runtime in days per pretraining experiment (40k iterations) for the different model configurations.
Model #Days GPUs
GPT-2 124M 2.2 4 AMD MI100
GPT-2 345M 8.8 4 AMD MI210
GPT-2 812M 14.1 4 AMD MI210
GPT-2 1.5B 20 4 AMD MI210
Table 8: Pretraining experiments runtime and GPU configurations used.
@@
A.2 Extended Analysis on Dropped and Retained Data Blocks

### v2

@@ 5.2 Baselines
    Industry-standard models: We compare the Llama-2 models trained through LFR with Pythia models Biderman et al. (2023) of similar scale obtained from EleutherAI’s Huggingface [^1] https://huggingface.co/models?other=pythia. These models have been trained on 300B tokens while the LFR models were trained on 9.8B tokens (3.2% of the tokens). We compare the GPT-2 models pretrained through LFR for 40k iterations with the same GPT-2 architectures pretrained by OpenAI  [^2] https://huggingface.co/openai-community for 800k iterations (5% of the training tokens). We use the same batch size as these industry-standard models (Section A.1) by adjusting the gradient accumulation steps and the per-device batch size.
@@
    Random Sampling: while the previous baselines ensures that we compare with industry-standard models, we also develop and compare LFR against the same models pretrained using random sampling with 5.10$\times$ and 20$\times$ more tokens than LFR for the Llama-2 and GPT-2 models respectively. This baseline enables LFR to produce higher quality models than those obtained through traditional autoregressive modeling when using much fewer tokens and training iterations.
@@
    Prior works: We compare our training methodology with the dataset generated through DSIR Xie et al. (2023) and MiniPile Kaddour (2023).
5.3 Performance on Downstream Tasks
We perform single-shot evaluations on common-sense reasoning tasks using Llama-2 models trained through LFR, comparing the accuracy norm with the baselines in Table 2. Note that the LFR models are trained on only 3.2% of the tokens used to train the Pythia models and 19.6% of the tokens used for the random-sampling baseline. Additionally, the Pythia models have between 1.3×1.3$\times$1.3 × and 2.0×2.0$\times$2.0 × more parameters. Despite this, LFR models outperform the Pythia models on 55.55% of the datasets and consistently surpass the random-sampling baseline.
We test the GPT-2 models on language modeling tasks and compare with the OpenAI baseline in Table 3 by measuring the PPL. Note that our models are trained on 5% of the training tokens as compared with the OpenAI models, further validating that data quality is more important than quantity. We find that the PPL reduction obtained by LFR increase as the dataset size increases (from WikiText-2 to 1BW). Also, smaller models show a larger $PPL$ reduction by using LFR than larger models. On average, using our approach, perplexity was reduced by 4.92, 3.26, 2.17, and 1.40 for the GPT-2 124M, 345M, 774M, and 1.5B models, respectively.
@@ 5.3 Performance on Downstream Tasks
We also test the LFR-trained models on standard benchmarks from the translation, question-answering, world knowledge, and problem solving domains in Table 4. LFR models trained with 20$\times$ fewer training iterations achieves better performance than models trained using random sampling. Details of each of the datasets is provided in Section A.1.
Model Iters WMT NQ MMLU
(BLEU) (Acc) STEM (Acc) HM (Acc) SS (Acc) Other (Acc) Avg. (Acc)
1.5B OpenAI 800k 11.5 4.1 24.5 24.8 24.0 27.8 25.3
1.5B LFR 40k 11.8 4.61 26.1 27.2 23.8 25.1 25.5
Table 4: LFR-trained GPT-2 models evaluated on translation (WMT-14 wmt ), question-answering (Natural Questions Kwiatkowski et al. (2019)), and world knowledge and problem solving (MMLU Hendrycks et al. (2021) domains using the BLEU scores and accuracy metrics. Note that NQ refers to Natural Questions, HM refers to Humanities, SS refers to Social Sciences, Other refers to business, health, and other miscellaneous topics, and Avg. refers to the average accuracy across all 57 subjects in MMLU. We compare our 1.5B parameter model with those trained by OpenAI for 20$\times$ more training iterations. The model with the superior performance is highlighted in bold.
@@ 5.4 Comparison with Prior Work
In this section, we use the 124M parameter model to compare LFR with DSIR Xie et al. (2023) and MiniPile Kaddour (2023), which propose importance resampling and cluster-based data sampling respectively. In the case of DSIR, we use the OpenWebText dataset as the raw dataset and set the target dataset as WikiText-103 for selecting a subset of 1M documents ($\sim$6GB). It took a total of 2.7 hours for the data selection process on 8 AMD EPYC 7V13 CPU cores operating at 2.45GHz with 64GB RAM. For the MiniPile baseline, we pretrain the GPT-2 124M model for 40k iterations with the dataset available on Huggingface (min, ).
Table 5 displays the downstream task perplexity. LFR significantly outperforms prior work on data selection and allows the pretrained LLM to generalize better to new finetuning objectives, by focusing and revising different parts of the training corpus at regular intervals.
Figure 4: PPL values are tracked at different training iterations for the clusters identified as challenging and prioritized during the Focus stage of LFR. The dotted line represents the PPL values for the same clusters when trained with random sampling (RS). Notably, LFR facilitates accelerated learning of these challenging data points between 20k and 60k iterations (the Focus stage), whereas random sampling consistently results in higher PPL values throughout.
Figure 5: PPL values are tracked at different training iterations for the clusters identified as easy, discarded during the Focus stage, and reintroduced during the Review phase. The dotted line represents the PPL values for the same clusters when trained with random sampling (RS). Notably, we demonstrate that models forget the data points discarded during training, unless reintroduced to the training corpus as in the case of LFR.
5.5 Ablation Study
In this section, our goal is to understand the impacts of the Focus and Review stages of LFR and exploring more aggressive data selection strategies by varying the hyperparameters $p_{1},s_{1},p_{2},p_{3},$ and $reps$.
5.5.1 Impact of Focus
Consider training the Llama-2 300M parameter model on the SlimPajama dataset, which comprises of seven sub-datasets sourced from CommonCrawl, Github, C4, Books, Wikipedia, StackExchange, and ArXiv. During the Focus stage, LFR employs weighted sampling from the three most challenging clusters while discarding clusters with the lowest perplexity (PPL). Additionally, within the retained clusters, LFR performs hierarchical sampling by prioritizing regions with higher PPL, further refining the data selection process. LFR classifies the Github, StackExchange, and ArXiv clusters as more challenging at 20k iterations, than the other four data sources. Figure 4 demonstrates the training dynamics of these challenging data points. Note that LFR (solid line) enables the accelerated learning of challenging parts of the dataset as opposed to random-sampling based training (dotted line).
Section A.2 provides examples of data considered challenging and easy from each of the sources in the SlimPajama dataset.
5.5.2 Impact of Review
Next, we study the impact of the Review phase on the data points identified as simple, and discarded during the Focus stage. Note that these samples are reintroduced to the training process in LFR, unlike prior work in data selection. Figure 5 demonstrates the importance of the Review phase by plotting the PPL values for the easy data points for LFR (solid line) and random sampling (dotted line). The model forgets data points which are discarded during training (solid line is higher than the dotted line in the Focus stage). This is because the model is focusing on more challenging clusters during this phase. Therefore, the Review phase is essential to achieving performant models, and provides LFR a distinct edge over other data selection methodologies evaluated in Section 5.4. Please refer to Section A.2 for examples of data samples identified as easy and difficult by LFR.
@@ 8 Limitations and Ethical Considerations
    LFR is evaluated on models up to 1.5B parameters using the OpenWebText dataset, constrained by our compute resources. With the clear success on models of such scale, we hope to inspire researchers to validate such focused learning approaches for different model families, and domains.
@@ A.1 Experiment Details
    ARC-Challenge arc : A subset of the AI2 Reasoning Challenge with 2,590 challenging multiple-choice science questions designed to test advanced reasoning and knowledge.
  2. 2.
    ARC-Easy arc : A subset of the AI2 Reasoning Challenge with 5,117 relatively easier multiple-choice science questions focusing on basic reasoning and recall.
  3. 3.
    BoolQ boo : A dataset of 16,000+ boolean (yes/no) questions paired with passages, requiring models to infer answers from supporting evidence.
  4. 4.
    HellaSwag hel : A dataset with 70,000+ multiple-choice questions focused on commonsense reasoning and contextual understanding, particularly in describing scenarios.
  5. 5.
    OpenBookQA Ope : A multiple-choice question-answering dataset with 5,957 questions requiring knowledge retrieval from a science "open book" and commonsense reasoning.
  6. 6.
    PIQA Piq : A physical commonsense reasoning dataset with 20,000+ binary-choice questions about everyday situations and physical interactions.
  7. 7.
    Winogrande win : A dataset with 44,000+ sentence pairs designed to test commonsense reasoning through pronoun disambiguation challenges.
  8. 8.
@@
Models: Tables 6 and 7 describes the different model configurations and pretraining hyperparameters used in LFR for the Llama-2 models.
120M 300M 500M
Layers 12 12 11
#Heads 12 16 32
n_embd 768 1024 2048
Table 6: Number of layers, attention heads, and the embedding dimensions in the Llama-2 models used for pretraining.
Parameter Value
Context Length 1024
Embedding Dimension (768, 1024, 2048)
Total Iterations 100,000
Effective Batch Size 768
Block Size 4096
Weight Decay 1.00E-1
Adam $\beta_{1}$ 0.90
Adam $\beta_{2}$ 0.95
Warmup Iterations 8000
Minimum Learning Rate 4.00E-5
Maximum Learning Rate 4.00E-04
Learning Rate Schedule Cosine Decay
Learning Rate Decay Iterations 100,000
GPUs (4x NVIDIA A6000)
Table 7: Pretraining hyperparameters for the Llama-2 120M-500M parameter models. Parameters with multiple values (e.g. Embedding dimensions, batch size, gradient accumulation steps, and GPUs) specified in brackets are for the 120M, 300M, and 500M parameter models respectively.
Tables 8 and 9 describes the different model configurations and pretraining hyperparameters used in LFR for the GPT-2 models.
@@
A.2 Llama Pretraining - Data Importance
Figure 6: Clustering the data embeddings from the SlimPajama dataset using the Llama-300M model at the 50k training step.
In this section, we study the data points identified as easy and challenging by LFR when pretraining with the SlimPajama dataset. Listing 1 provides an example of a code snippet from Github classified as easy by LFR, and discarded in the Focus stage of the Llama model training. Listing 2 provides an example of a data sample retained from the Github cluster. Note that this code is more complex, presents an opportunity to the model to improve its coding capabilities as opposed to the variable declarations in Listing 1.
Listing 1: Code snippet classified as easy by LFR, primarily consisting of variable declarations. As seen from the code, it contributes minimally to enhancing the model’s coding capabilities.
package frclibj;
import edu.wpi.first.wpilibj.Timer;
public class TrcDbgTrace
[…]
public static final String ESC_SUFFIX = "m";
public static final String ESC_SEP = ";";
public static final String SGR_RESET = "0";
public static final String SGR_BRIGHT = "1";
public static final String SGR_DIM = "2";
public static final String SGR_ITALIC = "3";
public static final String SGR_UNDERLINE = "4";
public static final String SGR_BLINKSLOW = "5";
public static final String SGR_BLINKFAST = "6";
public static final String SGR_REVERSE = "7";
public static final String SGR_HIDDEN = "8";
public static final String SGR_CROSSEDOUT = "9";
public static final String ESC_NORMAL = ESC_PREFIX;
Listing 2: Code snippet classified as challenging by LFR. This code consists of a function which executes an Oracle query and returns a scalar value. As seen from the code, it contributes significantly to enhancing the model’s coding capabilities as compared with Listing 1.
/// <summary>
/// Executes an Oracle query that returns a single scalar value as the result.
/// </summary>
/// <param name="commandText">The Oracle query to execute</param>
/// <param name="parameters">Optional parameters to pass to the query</param>
/// <returns>The result of the query as an object</returns>
public object QueryValue(string commandText, IEnumerable parameters)
object result;
if (String.IsNullOrEmpty(commandText))
throw new ArgumentException("Command text cannot be null or empty.");
try
ensureConnectionOpen();
var command = createCommand(commandText, parameters);
result = command.ExecuteScalar();
finally
ensureConnectionClosed();
return result;
Similarly, we also provide examples of question-answer pairs from StackExchange which were discarded and retained in the Focus stage of the Llama pretraining in Listings 3 and 4 respectively.
Listing 3: Question-answer pair from StackExchange classified as easy by LFR. The question revolves around a process in PayPal which does not contribute as much to the answering capability or world knowledge of the model.
Q: PayPal IPN \$_POST[’txn_id’] not set. Im using the PayPal sandbox to do a subscribe button, and then when I get the IPN response for a subscription or a subscription cancellation \$_POST[’txn_id’] is never set.
So I dont know how to identify transactions to only accept unique ones.
Thanks!
EDIT: for example, all the info that I have in POST for a subscr_cancel are:
amount1, amount3, address_status, subscr_date, payer_id, address_street, mc_amount1, mc_amount3, charset, address_zip, first_name, reattempt, address_country_code, address_name, notify_version, subscr_id, custom, payer_status, business, address_country, address_city, verify_sign, payer_email, btn_id, last_name, address_state, receiver_email, recurring, txn_type, item_name, mc_currency, residence_country, test_ipn, period1, period3, correlation_id.
A: According to Table 2. Summary of subscription variables:
For subscription variables, the transaction ID (txn_id) is only available for USD Payment and Multi-Currency Payment transaction types (txn_type).
As expected, PayPal will not send the txn_id to your IPN for the transaction type, subscr_cancel, and will only send txn_id if the transaction type is subscr_payment.
For further explanation on which variables are sent to your IPN URL based on your transaction, please check out IPN and PDT Variables.
Have you checked \$_REQUEST[’txn_id’] as this may be sent to your server via GET.
Listing 4: Question-answer pair from StackExchange classified as challenging by LFR. The question revolves around solving an ODE which contributes more to the learning of the model than Listing 3.
Q: Passing additional iteration-dependent inputs to ode45
Im trying to solve a differential equation using the ode45 function. Consider the following code,
[t1,X2] = ode45(@(t,x)fun(t,x,C1,C2,C3,C4),t0,X01);
where parameters C1, C2, C3, and C4 are column vectors, which should be available to the function that ode45 is referring to (fun.m).
I want the values to change after every iteration, so for example, at the beginning the entry of C1 I want is C1(1), in the next iteration its C1(2), etc.
How can I implement that?
A: You may have noticed that the official docs are not too helpful in this scenario (as they pretty much force you to use global variables - which is doable, but discouraged).
Instead, Ill show you how this can be done with classes and function handles. Consider the following:
classdef SimpleQueue < handle
%SIMPLEQUEUE A simple FIFO data structure.
properties (Access = private)
data
position
end
methods (Access = public)
function obj = SimpleQueue(inputData)
%SIMPLEQUEUE Construct an instance of this class
obj.data = inputData;
rewind(obj);
end % constructor
function out = pop(obj, howMany)
%POP return the next howMany elements.
if nargin < 2
howMany = 1; % default amount of values to return
end
finalPosition = obj.position + howMany;
if finalPosition > numel(obj.data)
error(’Too many elements requested!’);
end
out = obj.data(obj.position + 1 : obj.position + howMany);
obj.position = finalPosition;
end % pop
function [] = rewind(obj)
%REWIND restarts the element tracking
% Subsequent calls to pop() shall return elements from the beginning.
obj.position = 0;
end % rewind
end % methods
end % classdef
How to use this? Simple:
C1q = SimpleQueue(C1);
C2q = SimpleQueue(C2);
C3q = SimpleQueue(C3);
C4q = SimpleQueue(C4);
[t1,X2] = ode45(@(t,x)fun(t,x,@C1q.pop,@C2q.pop,
@C3q.pop,@C4q.pop),t0,X01);
As you can see, inside fun we use C1q() instead of C1.
A.3 Sensitivity Study
In this section, our goal is to understand the effects of more aggressive focus, revision, and learning strategies than the training strategy presented in Section 4. Here, we vary the values of hyperparameters $p_{1},s_{1},p_{2},p_{3},$ and $reps$ and study the effects on the downstream task perplexity. Note that the GPT-2 models used a four phase training process. Specifically, we aim to answer the following two questions using the GPT-2 models:
  1. 1.
    What is the impact of not reintroducing the discarded data samples?
  2. 2.
    What is the impact of the degree of pruning in Phases 2 and 4?
To answer the first question, we pretrain a 124M parameter GPT-2 model without the reintroduction of data blocks in Phase 3, and use the reduced subset from Phase 2 for the rest of the training. Then, we finetune for downstream language modeling tasks similarly and compared the perplexities using LFR in Table 10. This training strategy which removes Phase 3, is labeled as no-reintro. Next, to answer the second question, we pretrain a 124M parameter GPT-2 model using LFR but increase the degree of pruning in Phase 2 from 50% to 70%, i.e., reduce the training subset to 30% of the original size. This aggressive training strategy is labeled as aggr-2.
A.4 Analysis on Dropped and Retained Data Blocks - GPT-2
  1. 1.
  2. 2.
  3. 3.
    Are the dropped data blocks similar across model sizes?
  4. 4.
To answer the first question, we printed out the texts dropped and retained at different training phases. Tables 14 and 16 show text blocks dropped in Phases 2 and 4 by the 345M and 124M parameter models, while Tables 15 and 17 show data blocks retained. By reading through the texts, we notice that the model first learned conversations and personal anecdotes, before being able to retain factual information. We provide a more detailed analysis of the learning process in Section A.5.
To summarize, data block importance varies across training time, and across model sizes. Therefore, static data selection techniques Tirumala et al. (2023); Abbas et al. (2023); Kaddour (2023); Xie et al. (2023) which select a fixed subset to train for the entire training duration for all model architectures do not adapt to the changing training dynamics of LLMs. Based on our analysis in Figure 8 and 7, different data blocks are found difficult by models of different capacities at different training instants, driving the need for dynamic data selection methods like LFR. We detail further analysis on the selected and discarded data blocks and demonstrate how models initially focus on learning conversational and anecdotal data, before proceeding to learn factual data in Appendix A.5.
A.5 Extended Analysis on Dropped and Retained Data Blocks for GPT-2